OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 3) and (2, 7)? ANSWERS BELOW PLS HELP!

OpenStudy (anonymous):

x - y = -1 x - y = -2 x - y = - 5 x - y = - 6